A stone is dropped from the top of the tower of height 60m.Simultaneously another stone is projected vertically upwards from the foot of the tower .They meet at a height 2h/3 from the ground level .The initial velocity of the stone projected upwards is
 

Grade:11

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Time by upper stone to cover h/3 = 20 m is 2 sec

Time by lower stone to cover 40 m = 2 sec

Finding initial velocity

40 = 2u - 0.5*10*2*2
60/ 2 = 30 m/s = u

u = 30 m/s

Regards